Successive discount of a percent and bpercent are equivalent to a single discount of ?

Grade:8

2 Answers

Arun
5 years ago
Dear student
 
You have forgot to insert the discounted value. Please check and repost this.
 
 
You can also attach an image.
 
I will be happy to help you
Sayantan Garai
5 years ago
 
If M is the marked price then the selling price is
 (1-A/100)(1-B/100)×M.
Thus if M=100,
SP=(1-A/100)(1-B/100)×100.
Hence, Discount=(M-SP)=100-SP
             => D(%)= (M-SP)/M ×100= 100-SP.